# thanks: Corresponding author

### v3

@@ I Introduction
VDM is described by a four vector field $A_{\mu}$. Given a mass, $m$, this leads to three polarization states: two transverse helicities, $A^{\pm}_{T}$, related to circular polarization of light, and one longitudinal mode, $A_{L}$, related to the mass. Most production mechanisms for VDM Graham et al. (2016); Agrawal et al. (2020); Dror et al. (2019); Long and Wang (2019) lead to either the longitudinal mode, or the transverse mode, or both, being populated. As described in detail below and in Refs. Zhang et al. (2022); Amin et al. (2022); Gorghetto et al. (2022); Adshead and Lozanov (2021); Jain and Amin (2022); Gosenca et al. (2023); Glennon et al. (2023); Chen et al. (2023), in the non-relativistic limit appropriate for DM in our galaxy VDM is described by three Schrödinger fields $\psi_{i}$ aligned with the Cartesian coordinate axes and composed into a vector $\bm{\Psi}$, with the circular/elliptical polarization now encoded in the spin angular momentum density $\bm{s}=i\bm{\Psi}\times\bm{\Psi}^{*}$. Initial conditions of either longitudinal or transverse polarizations in terms of $A_{\mu}$ in general populate all $\psi_{i}$ modes, and thus give rise a random distribution for the spin polarization (see e.g. Ref. Gorghetto et al. (2022); Jain et al. (2023a); Amaral et al. (2024)). Even in such a random model, however, the polarization will (just like scalar/axion DM Marsh (2016); O’Hare (2024)) display a coherence length and time governed by gravitational dynamics. On the other hand, there also exist Bianchi universe and Lorentz violating models that can preferentially populate given $\psi_{i}$ modes and give rise to directionally polarized states Arias et al. (2012); Alonso-Álvarez et al. (2020) (although see Refs. Himmetoglu et al. (2009a, b); Karciauskas and Lyth (2010)).
@@
Ref. Caputo et al. (2021) noted that, at the time of publication, the precise spin-vector distributions after accounting for non-linear structure formation had not been considered. It is the goal of the present work to study exactly this effect. Previous works have studied the collapse and condensation of vector solitons and DM halos Chen et al. (2023); Jain et al. (2023a); Gorghetto et al. (2022); Amin et al. (2022). It has been found that non-linear gravitational collapse leads to a kind of thermalisation and equipartition between the spin polarization modes Amaral et al. (2024). Ref. Amaral et al. (2024) considered how the polarization evolves on timescales of order the coherence time, $\tau_{c}=1/mv_{\rm vir}^{2}$ with $v_{\rm vir}$ the halo virial velocity, finding that the polarization follows the density coherence and that the vector field sweeps out ellipses in real space.
@@
We consider a series of different simulations with different initial spin states for VDM. Our simulations are not cosmological, but instead they condense VDM halos from homogeneous an isotropic noise, following Refs. Levkov et al. (2018); Chen et al. (2021, 2023); Jain et al. (2023a); Chen (2021). The statistical properties of the DM in such halos are known to be similar to those formed cosmologically (e.g. Ref. Schive et al. (2014a, b)), but can be simulated at lower computational cost. We consider models with random spin polarization, i.e. random mixes between transverse and longitudinal polarizations. We also consider, for the first time, a novel mixed state polarization which mixes the fixed and random models. We do not simulate fully polarized cases (either spin or linear), as these have trivial evolution in terms of polarization.
@@ II Equations of Motion and Initial Conditions
where $\bm{\Psi}=(\psi_{1},\psi_{2},\psi_{3})^{T}$ is the wave function vector, $n=n_{1}+n_{2}+n_{3}$ is the mean total number density, $G$ is Newton’s gravitational constant, and natural units, $\hbar=c=1$, are used. The number density, mass density, and spin density are $N=\bm{\Psi}^{*}\bm{\Psi}$, $\rho=m\bm{\Psi}^{*}\bm{\Psi}$, $\bm{s}=\left(i\bm{\Psi}\times\bm{\Psi}^{*}\right)$. The $SO(3)$ symmetry leads to the conservation of intrinsic spin angular momentum, $\bm{S}$:
@@
Similarly to previous studies on the condensation of VDM Chen et al. (2023); Jain et al. (2023a), we generate homogeneous and isotropic initial conditions in a periodic box of size $L$. Each component of the vector field is assumed to be a Gaussian random field with a momentum distribution of $|\psi_{i}(\bm{p})|^{2}\propto\delta(|\bm{p}|-mv_{0})$. Specifically, given the momentum distribution we perform an inverse Fourier transform on $\psi_{i}(\bm{p})e^{iA}$ with $A$ a random phase uniformly drawn from $[0,2\pi)$ for each non-zero $\psi_{i}(\bm{p})$. The wave function is normalized to give a specified total number of non-relativistic bosons for the $i-$th component in the box, $N_{i}\equiv n_{i}L^{3}$. The initial condition is then evolved by solving the SP equations numerically using a fourth-order time-splitting pseudospectral method Du et al. (2018); Mocz et al. (2017); Chen et al. (2023, 2021, 2022); Chen and Zhang (2024).
Figure 1: Slice of particle number density (left figure) and spin density (right figure) evolution for the three initial conditions within the simulation box of $\widetilde{L}=64$. In each figure, top row indicates the formation of a halo and Proca star from homogeneous and isotropic CF initial conditions with $\widetilde{N}_{1}=\widetilde{N}_{2}=\widetilde{N}_{3}\approx 630$. Middle row is the evolution for the RF initial condition. Bottom row is for CF/$\psi_{i}=0$ initial condition . Here we take $\widetilde{\psi_{x}}=\psi_{1}$, $\widetilde{\psi_{y}}=(2\psi_{2}-i\psi_{1})/\sqrt{5}$, $\widetilde{\psi}_{z}=0$ with $\widetilde{N}_{1}=\widetilde{N}_{2}\approx 630$, $\widetilde{N}_{3}=0$. The figures on the right, from top to bottom, respectively refer to the slice of spin density.
@@
    CF – three independent complex fields. In this case, a random polarization value occurs throughout the box at the initial time.
@@
    RF – three independent real fields. In this case, the polarization is exactly zero everywhere at the initial time by symmetry.
@@
    Correlated initial conditions, specified by the correlation matrix $C_{ij}$ Chen et al. (2023). Taking $\psi_{z}=0$ with different degree of correlation between $\psi_{x}$ and $\psi_{y}$, this state leads to a partially polarized initial condition with non-zero $\langle S^{z}\rangle$.
@@ III.1 Global Properties of Density and Spin
Fig.1 shows slices through our simulation boxes for each initial condition, displaying the VDM density and spin. The density is shown in the left column of Fig.1, where we observe the formation of a halo with a Proca star in its center. Due to the identical total mass conditions in CF and RF, the time for the formation of stable halos and Proca stars is roughly the same, occurring around $\widetilde{t}=12000$. In the case of CF/$\psi_{i}=0$, however, due to the smaller total mass and the correlation among the components, halos and Proca stars had already formed at $\widetilde{t}=4500$ and $\widetilde{t}=7500$ respectively. This is consistent with the results in Ref. Chen et al. (2023).
@@ III.2 Uncorrelated initial conditions
Despite zero spin density on average over most of the simulation, the Proca star nonetheless develops polarization. This is demonstrated in Fig.4 which shows $\widetilde{S}^{i}/\widetilde{N}$ evaluated for the Proca star. After formation, the norm of the spin develops a non-zero value, which grows and undergoes strong oscillations with time. The Proca star has $\widetilde{S}^{i}/\widetilde{N}<1$ and appears to approach the limiting case $\widetilde{S}^{i}/\widetilde{N}=1$ as time goes on. The reason is that in the case of gravity alone, different spins of solitons can overlap, and the resultant is still a soliton, depending on whether there are more directional solitons, average total spin of 0, or more spinning solitons, average total spin of 1, then the soliton formed by the superposition can take any value between 0 and 1.
@@ III.3 Spin polarization from correlated initial conditions
Compared to the previous unpolarized state where each component $\psi_{i}$ of the system was independent in the initial condition, we now set $\psi_{z}=0$ and correlate $\psi_{x,y}$ using Eqs. (10, 11). This gives a partially circularly polarized state initially. By employing Equation Eq. 6, it is straightforward to deduce that under these conditions only the $S^{z}$ component possesses spin. With a high degree of correlation, we enforce a large and resolvable (larger than numerical error) value to $S^{z}$. Symmetry dictates that the other components of the spin remain exactly zero everywhere, which the system respects. The spin in the halo is shown in Fig.7, while the spin of the core is shown in Fig.8.
@@
Figure 7: Correlated initial conditions. The three sub-figures above are the evolution of average spin density and its error bar in different area, $\geq 0.2,0.5,1.0\widetilde{r}_{vir}$. The bottom subfigure shows the growth of maximum density. The blue line is $\widetilde{t}=4500$ and $\widetilde{t}=7500$, representing the formation of halo and proca star, respectively.Here $\widetilde{r}_{vir}\approx 24.25$, initial $\widetilde{S}^{z}/\widetilde{N}=-0.456$, $\widetilde{S}^{x}/\widetilde{N}=\widetilde{S}^{y}/\widetilde{N}=0$
@@ IV Spin Mixed State of VDM
    $\kappa=1$: The VDM population is composed of spin-aligned VDM particles, meaning at least one average spin component is non-zero, and the average spin amplitude is equal to one ($\widetilde{S}/\widetilde{N}=1$), as depicted in the bottom-right sub-figure of Fig.9. The environment where $\kappa=1$ can exist in the core region where the spin density asymptotically increase to one as the simulation time reaches very large values (see Fig.8). VDM could also have an initial condition where the correlation is highest, resulting in an initial spin density equal to one, e.g., $C_{1}=1/\sqrt{2},C_{2}=i/\sqrt{2}$ Chen et al. (2023).
@@ V Implication to Direct Searches
Experiments that directly search for VDM, called haloscopes, utilize the conversion of VDM particles to photons within the geometry of the experiment, utilizing the kinetic mixing coupling between the detectable electric field and the dark (or “hidden”) electric field of the VDM. Thus, in this section, we focus on the behavior of the detectable signal power of VDM within the haloscope. The interpretation is based on the dynamics of VDM in the outer halo region (Sec. III, in a location that is geometrically comparable to the location of the Solar System in the Milky Way. Starting from the properties of the spin of VDM particles, one can obtain the behaviour of the dark electric field of VDM given by the following equation Amaral et al. (2024):
@@
Thus, non-zero spin defines circular polarization of the dark electric field $\bm{A}$. The rotation period of the polarization is given by the Compton period of the VDM ($T_{\mathrm{C}}=1/m_{\mathrm{VDM}}$). Circular polarization (as emphasised by e.g. Refs. Amin et al. (2022); Amaral et al. (2024)) presents a stark contrast with the behavior of $\bm{A}$ widely assumed by the VDM haloscope search community, which is linear polarization i.e., $\bm{A}$ of a VDM particle oscillates in one direction, so $\bm{S}=0$. The community generally assumes linear polarization for $\bm{A}$, where the random polarization case is equivalent to the random distribution of the direction of the dark electric field, while fixed polarization over long time scales (much larger than sidereal period) corresponds to the aligned direction of the dark electric field Arias et al. (2012). Ref. Caputo et al. (2021) discusses the sidereal modulation of the VDM signal in the latter case of fixed polarization. This linear polarization condition can only exist under very specific conditions of the VDM, which result in $\widetilde{S}=0$ and differs from our findings of non-zero $\bm{S}$ for each VDM particle.
@@
Figure 10: Distribution of the dark electric field $\bm{A}_{\mathrm{DM}}$ under the spin’s mixed state. The cyan circles denote the random spin population of VDM, which results in randomly distributed $\bm{A}_{\mathrm{DM}}$. The black circles denote the aligned spin population, resulting in $\bm{A}_{\mathrm{DM}}$ collectively pointing in the same direction and coherently rotating. The $\bm{A}_{\mathrm{DM}}$ of each VDM particle has a frequency of rotation determined by its mass.
@@
To demonstrate with a case study, we choose the BRASS-p experiment Bajjali et al. (2023) due to its world-leading sensitivity for VDM within the mass range of (41.35 - 74.44) $\mu$eV. BRASS-p is equipped with a state-of-the-art radio telescope’s full-fledged analog and digital backend, allowing it to measure the electric field using its double linear polarization receiver. As a result, the signal recovered is not only equal to the power detected by the square-law detector, but the GPU-powered digital backend can also dissect the incoming signal to obtain the four Stokes parameters of the polarization signalRobishaw and Heiles . In this case study, the chosen spin-vector of the aligned spin population is $\hat{\bm{S}}=(1/4,\sqrt{3}/4,\sqrt{3}/2)^{T}$, $\hat{\bm{A}}(t=0)$ can be chosen arbitrarily and $\hat{\bm{S}}\perp\hat{\bm{A}}$. Time evolution of $\hat{\bm{A}}$ in Compton period can be obtain by the Rodrigues’ rotation formula about the vector $\hat{\bm{S}}$
@@
Fig. 10 shows the sidereal modulation of $(\hat{\bm{A}}\cdot\hat{\bm{G}})^{2}$ for the aligned spin VDM in the sidereal and Compton time scales. The polarization rotation yields a conversion rate that oscillates at the Compton time scale between zero (when $\hat{\bm{A}}\perp\hat{\bm{G}}$) and the upper value given by the configuration of $\hat{\bm{G}}$ in the sidereal time scale. Haloscope experiments typically involve long data-taking campaigns composed of many short periods ($\gg T_{\mathrm{C}}$), which are combined to produce a single spectrum. As a result, signal modulation on the Compton time scale is averaged out in the average spectrum. The averaging of the $(\hat{\bm{A}}\cdot\hat{\bm{G}})^{2}$ within the Compton period is a suitable way to present the signal per instance of data-taking, and thus the signal in the average spectrum is only subjected to the sidereal modulation. Because of this averaging in the Compton time scale, the modulation span appears to be milder than the linear polarization discussed in Caputo et al. (2021).

### v4

@@ I Introduction
VDM is described by a four vector field $A_{\mu}$. Given a mass, $m$, this leads to three polarization states: two transverse helicities, $A^{\pm}_{T}$, related to circular polarization of light, and one longitudinal mode, $A_{L}$, related to the mass. Most production mechanisms for VDM Graham et al. (2016); Agrawal et al. (2020); Dror et al. (2019); Long and Wang (2019) lead to either the longitudinal mode, or the transverse mode, or both, being populated. As described in detail below and in Refs. Zhang et al. (2022); Amin et al. (2022); Gorghetto et al. (2022); Adshead and Lozanov (2021); Jain and Amin (2022); Gosenca et al. (2023); Glennon et al. (2023); Chen et al. (2023a), in the non-relativistic limit appropriate for DM in our galaxy VDM is described by three Schrödinger fields $\psi_{i}$ aligned with the Cartesian coordinate axes and composed into a vector $\bm{\Psi}$. The fields $\psi_{i}$ give the amplitude of the vector potential $A_{i}$ in the non-relativistic limit (WKB approximation) in which the “fast” Compton timescale, $1/m$ is factored out. The $\psi_{i}$ fields interact with one another via a non-linear Schrödinger equation, with the potential determined via Poisson’s equation for the mass density of VDM, which is proportional to $|\bm{\Psi}|^{2}$. This is in exact analogy with the case of scalar/axion DM (see e.g. Ref. Marsh (2015) for a derivation of the Schrödinger-Poisson equation from the relativistic Klein-Gordon-Einstein equations). It is important to note that the relevant description arises from the classical field equations: the appearance of the Schrödinger equation is a mathematical coincidence and is not the quantum equation of the same name.
Circular/elliptical polarization is now encoded in the spin angular momentum density $\bm{s}=i\bm{\Psi}\times\bm{\Psi}^{*}$. Initial conditions of either longitudinal or transverse polarizations in terms of $A_{\mu}$ in general populate all $\psi_{i}$ modes, and thus give rise a random distribution for the spin polarization (see e.g. Ref. Gorghetto et al. (2022); Jain et al. (2023a); Amaral et al. (2024)). Even in such a random model, however, the polarization will (just like scalar/axion DM Marsh (2016); O’Hare (2024)) display a coherence length and time governed by gravitational dynamics. On the other hand, there also exist Bianchi universe and Lorentz violating models that can preferentially populate given $\psi_{i}$ modes and give rise to directionally polarized states Arias et al. (2012); Alonso-Álvarez et al. (2020) (although see Refs. Himmetoglu et al. (2009a, b); Karciauskas and Lyth (2010)).
@@
Ref. Caputo et al. (2021) noted that, at the time of publication, the precise spin-vector distributions after accounting for non-linear structure formation had not been considered. It is the goal of the present work to study exactly this effect. Previous works have studied the collapse and condensation of vector solitons and DM halos Chen et al. (2023a); Jain et al. (2023a); Gorghetto et al. (2022); Amin et al. (2022). It has been found that non-linear gravitational collapse leads to a kind of thermalisation and equipartition between the spin polarization modes Amaral et al. (2024). Ref. Amaral et al. (2024) considered how the polarization evolves on timescales of order the coherence time, $\tau_{c}=1/mv_{\rm vir}^{2}$ with $v_{\rm vir}$ the halo virial velocity, finding that the polarization follows the density coherence and that the vector field sweeps out ellipses in real space.
@@
We consider a series of different simulations with different initial spin states for VDM. Our simulations are not cosmological, but instead they condense VDM halos from homogeneous an isotropic noise, following Refs. Levkov et al. (2018); Chen et al. (2021, 2023a); Jain et al. (2023a); Chen (2021). The statistical properties of the DM in such halos are known to be similar to those formed cosmologically (e.g. Ref. Schive et al. (2014a, b)), but can be simulated at lower computational cost. We consider models with random spin polarization, i.e. random mixes between transverse and longitudinal polarizations. We also consider, for the first time, a novel mixed state polarization which mixes the fixed and random models. We do not simulate fully polarized cases (either spin or linear), as these have trivial evolution in terms of polarization.
@@ II Equations of Motion and Initial Conditions
where $\bm{\Psi}=(\psi_{1},\psi_{2},\psi_{3})^{T}$ is the wave function vector, $n=n_{1}+n_{2}+n_{3}$ is the mean total number density, $G$ is Newton’s gravitational constant, and natural units, $\hbar=c=1$, are used. The number density, mass density, and spin density are $N=\bm{\Psi}^{*}\bm{\Psi}$, $\rho=m\bm{\Psi}^{*}\bm{\Psi}$, $\bm{s}=\left(i\bm{\Psi}\times\bm{\Psi}^{*}\right)$. The spin density $\bm{s}$ arises from the $\mathrm{SO}(3)$ symmetry of $\bm{\Psi}=(\psi_{1},\psi_{2},\psi_{3})^{T}$, describing the spatial distribution of the conserved angular momentum. The $SO(3)$ symmetry leads to the conservation of intrinsic spin angular momentum, $\bm{S}$:
@@
Similarly to previous studies on the condensation of VDM Chen et al. (2023a); Jain et al. (2023a), we generate homogeneous and isotropic initial conditions in a periodic box of size $L$. Each component of the vector field is assumed to be a Gaussian random field with a momentum distribution of $|\psi_{i}(\bm{p})|^{2}\propto\delta(|\bm{p}|-mv_{0})$. Specifically, given the momentum distribution we perform an inverse Fourier transform on $\psi_{i}(\bm{p})e^{iA}$ with $A$ a random phase uniformly drawn from $[0,2\pi)$ for each non-zero $\psi_{i}(\bm{p})$. The wave function is normalized to give a specified total number of non-relativistic bosons for the $i-$th component in the box, $N_{i}\equiv n_{i}L^{3}$. The initial condition is then evolved by solving the SP equations numerically using a fourth-order time-splitting pseudospectral method Du et al. (2018); Mocz et al. (2017); Chen et al. (2023a, 2021, 2022); Chen and Zhang (2024).
Figure 1: Slice of particle number density (left figure) and spin density (right figure) evolution for the three initial conditions within the simulation box of $\widetilde{L}=64$. In each figure, top row indicates the formation of a halo and Proca star from homogeneous and isotropic CF initial conditions with $\widetilde{N}_{1}=\widetilde{N}_{2}=\widetilde{N}_{3}\approx 630$. The middle row is the evolution for the RF initial condition (exact zero spin density). The bottom row is for the CF/$\psi_{i}=0$ correlated initial condition with partial polarization. Here we take $\widetilde{\psi_{x}}=\psi_{1}$, $\widetilde{\psi_{y}}=(2\psi_{2}-i\psi_{1})/\sqrt{5}$, $\widetilde{\psi}_{z}=0$ with $\widetilde{N}_{1}=\widetilde{N}_{2}\approx 630$, $\widetilde{N}_{3}=0$. The figures on the right, from top to bottom, respectively refer to the slice of spin density.
@@
    CF – three independent complex fields. We first use the initial conditions employed in the study of Proca star condensation time Chen et al. (2023a). In this case, a random polarization value occurs throughout the box at the initial time.
@@
    RF – three independent real fields. This initial condition has zero spin density, and maintains it exactly due to symmetry. It serves as a test of our numerical and physical understanding.
@@
    Correlated initial conditions, (labeled CF/$\psi_{i}=0$). In addition to the initial conditions with independent components, we also consider those specified by the correlation matrix $C_{ij}$ Chen et al. (2023a). Taking $\psi_{z}=0$ with different degrees of correlation between $\psi_{x}$ and $\psi_{y}$, this state leads to a novel partially polarized initial condition with non-zero $\langle S^{z}\rangle$.
@@ III.1 Global Properties of Density and Spin
Fig.1 shows slices through our simulation boxes for each initial condition, displaying the VDM density and spin. The density is shown in the left column of Fig.1, where we observe the formation of a halo with a Proca star in its center. Due to the identical total mass conditions in CF and RF, the time for the formation of stable halos and Proca stars is roughly the same, occurring around $\widetilde{t}=12000$. In the case of CF/$\psi_{i}=0$, however, due to the smaller total mass and the correlation among the components, halos and Proca stars had already formed at $\widetilde{t}=4500$ and $\widetilde{t}=7500$ respectively. This is consistent with the results in Ref. Chen et al. (2023a).
@@ III.2 Uncorrelated initial conditions
Despite zero spin density on average over most of the simulation, the Proca star nonetheless develops polarization. This is demonstrated in Fig.4 which shows $\widetilde{S}^{i}/\widetilde{N}$ evaluated for the Proca star, we define the radius of Proca star, $r_{\mathrm{core}}=\max(\mathrm{density})^{-0.25}\times 1.08$ Schive et al. (2014a); Chen et al. (2023b). After formation, the norm of the spin develops a non-zero value, which grows and undergoes strong oscillations with time. The Proca star has $\widetilde{S}^{i}/\widetilde{N}<1$ and appears to approach the limiting case $\widetilde{S}^{i}/\widetilde{N}=1$ as time goes on. The reason is that in the case of gravity alone, different spins of solitons can overlap, and the resultant is still a soliton, depending on whether there are more directional solitons, average total spin of 0, or more spinning solitons, average total spin of 1, then the soliton formed by the superposition can take any value between 0 and 1.
@@ III.3 Spin polarization from correlated initial conditions
Compared to the previous unpolarized state where each component $\psi_{i}$ of the system was independent in the initial condition, we now set $\psi_{z}=0$ and correlate $\psi_{x,y}$ using Eqs. (10, 11). This gives a partially circularly polarized state initially. By employing Eq. 6, it is straightforward to deduce that under these conditions only the $S^{z}$ component possesses spin. With a high degree of correlation, we enforce a large and resolvable (larger than numerical error) value to $S^{z}$. Symmetry dictates that the other components of the spin remain exactly zero everywhere, which the system respects. The spin in the halo is shown in Fig.7, while the spin of the core is shown in Fig.8.
@@
Figure 7: Correlated initial conditions. The three sub-figures above are the evolution of average spin density and its error bar in different area, $\geq 0.2,0.5,1.0\widetilde{r}_{vir}$. The bottom subfigure shows the growth of maximum density. The blue line is $\widetilde{t}=4500$ and $\widetilde{t}=7500$, representing the formation of halo and proca star, respectively.Here $\widetilde{r}_{vir}\approx 24.25$, initial $\widetilde{S}^{z}/\widetilde{N}=-0.456$, $\widetilde{S}^{x}/\widetilde{N}=\widetilde{S}^{y}/\widetilde{N}=0$.
@@ IV Spin Mixed State of VDM
    $\kappa=1$: The VDM population is composed of spin-aligned VDM particles, meaning at least one average spin component is non-zero, and the average spin amplitude is equal to one ($\widetilde{S}/\widetilde{N}=1$), as depicted in the bottom-right sub-figure of Fig.9. The environment where $\kappa=1$ can exist in the core region where the spin density asymptotically increase to one as the simulation time reaches very large values (see Fig.8). VDM could also have an initial condition where the correlation is highest, resulting in an initial spin density equal to one, e.g., $C_{1}=1/\sqrt{2},C_{2}=i/\sqrt{2}$ Chen et al. (2023a).
@@ V Implication to Direct Searches
Experiments that directly search for VDM, called haloscopes, utilize the conversion of VDM particles to photons within the geometry of the experiment, utilizing the kinetic mixing coupling between the detectable electric field and the dark (or “hidden”) electric field of the VDM. Thus, in this section, we focus on the behavior of the detectable signal power of VDM within the haloscope. The interpretation is based on the dynamics of VDM in the outer halo region (Sec. III, in a location that is geometrically comparable to the location of the Solar System in the Milky Way. Starting from the properties of the spin of VDM particles, one can obtain the behaviour of the dark electric field, $\bm{E}\approx\partial_{t}\bm{A}$, of VDM given by the following equation Amaral et al. (2024):
@@
Thus, non-zero spin defines circular polarization of the dark electric field. The rotation period of the polarization is given by the Compton period of the VDM ($T_{\mathrm{C}}=1/m_{\mathrm{VDM}}$). This time scale is also the period of the detectable EM signal in the haloscope experiment. Circular polarization (as emphasised by e.g. Refs. Amin et al. (2022); Amaral et al. (2024)) presents a stark contrast with the behavior of $\bm{A}$ widely assumed by the VDM haloscope search community, which is linear polarization i.e., $\bm{A}$ of a VDM particle oscillates in one direction, so $\bm{S}=0$. The community generally assumes linear polarization for $\bm{A}$, where the random polarization case is equivalent to the random distribution of the direction of the dark electric field, while fixed polarization over long time scales (much larger than sidereal period) corresponds to the aligned direction of the dark electric field Arias et al. (2012). Ref. Caputo et al. (2021) discusses the sidereal modulation of the VDM signal in the latter case of fixed polarization. This linear polarization condition can only exist under very specific conditions of the VDM, which result in $\widetilde{S}=0$ and differs from our findings of non-zero $\bm{S}$ for each VDM particle.
@@
Figure 10: Distribution of the dark electric field $\bm{A}$ under the spin’s mixed state. The cyan circles denote the random spin population of VDM, which results in randomly distributed $\bm{A}$. The black circles denote the aligned spin population, resulting in $\bm{A}$ collectively pointing in the same direction and coherently rotating. The $\bm{A}$ of each VDM particle has a frequency of rotation determined by its mass.
@@
To demonstrate with a case study, we choose the BRASS-p experiment Bajjali et al. (2023) due to its world-leading sensitivity for VDM within the mass range of (41.35 - 74.44) $\mu$eV. BRASS-p is equipped with a state-of-the-art radio telescope’s full-fledged analog and digital backend, allowing it to measure the electric field using its double linear polarization receiver. As a result, the signal recovered is not only equal to the power detected by the square-law detector, but the GPU-powered digital backend can also dissect the incoming signal to obtain the four Stokes parameters of the polarization signalRobishaw and Heiles . In this case study, the chosen spin-vector of the aligned spin population is $\hat{\bm{S}}=(1/4,\sqrt{3}/4,\sqrt{3}/2)^{T}$, $\hat{\bm{A}}(t=0)$ can be chosen arbitrarily and $\hat{\bm{S}}\perp\hat{\bm{A}}$. Time evolution of $\hat{\bm{A}}$ in Compton period can be obtain by the Rodrigues’ rotation formula around the vector $\hat{\bm{S}}$
@@
Fig. 10 shows the sidereal modulation of $(\hat{\bm{A}}\cdot\hat{\bm{G}})^{2}$ for the aligned spin VDM in the sidereal and Compton time scales. The polarization rotation yields a conversion rate that oscillates at the Compton time scale between zero (when $\hat{\bm{A}}\perp\hat{\bm{G}}$) and the upper value given by the configuration of $\hat{\bm{G}}$ in the sidereal time scale. Haloscope experiments typically involve long data-taking campaigns composed of many short periods ($\gg T_{\mathrm{C}}$), which are combined to produce a single spectrum. As a result, signal modulation on the Compton time scale is averaged out in the average spectrum. The averaging of the $(\hat{\bm{A}}\cdot\hat{\bm{G}})^{2}$ within the Compton period is a suitable way to present the signal per instance of data-taking, and thus the signal in the average spectrum is only subjected to the sidereal modulation. Because of this averaging in the Compton time scale, the modulation span appears to be milder than the linear polarization discussed in Caputo et al. (2021). As a side note, in haloscope experiments sensitive to very low-mass VDM where the coherence time $T_{\mathrm{C}}$ is comparable to or greater than the observation time $T_{\mathrm{obs}}$ or the sidereal time, the signal power depends on both $T_{\mathrm{C}}$ and $T_{\mathrm{obs}}$.